Three times one number added to five times another number is 54. The second number is two less than the first. Find the numbers.
Answer by checkley71(8403)

3X+5Y=54 & 5Y=3X-2 THUS 3X=54-5Y THUS 5Y=(54-5Y)-2 OR 5Y=54-5Y-2 OR 10Y=52 OR Y=52/10 OR Y=5.2 THEN 5*5.2=3X-2 OR 26=3X-2 OR 3X=28 OR X=28/3